# Imaging-spectroscopy diagnosis of the giant sloshing spiral in the Virgo cluster with the Einstein Probe Follow-up X-ray Telescope

### v2

@@ 4 Imaging-spectral analysis
The X-ray image obtained with EP-FXT was divided into 142 regions using the contour-binning software (Sanders 2006) with a SNR of 300. We adopted a flat-weighting assumption, assuming that the X-ray emission within each region is spatially uniform, for the Ancillary Response Files (ARFs) calculation. The ARFs were generated using the EP-FXT response tool fxtarfgen. During the calculation, the parameter extend was set to 1 to account for the extended nature of the source, thus more accurately capturing the spatial distribution of the emission within each extraction region. The spectra were grouped using grppha with a minimum of 10 counts per bin to ensure the applicability of the C-statistic. To minimize the impact of background contamination on the source spectra and improve the accuracy of the spectral analysis, we selected observations of the M87 source and its corresponding background, ensuring similar observation times, the same filter settings, and longer exposure durations; specifically, ObsID 13600005124 and ObsID 13600005125. In the spectral analysis, both the particle background and the diffuse background were taken into account. We calculated the particle background for both observations, with a ratio of $1.05\pm 0.03$, indicating that the particle background levels are nearly identical. Therefore, we selected the spectra from the same CCD region in ObsID 13600005125 to serve as the background for ObsID 13600005124. For regions in ObsID 13600005125 that are contaminated by bright point sources or extended sources, background data from nearby uncontaminated areas were used instead. Given that our analysis focuses on M87, we intentionally retained the Virgo cluster component in the background data because it more accurately reflects the diffuse emission environment surrounding the galaxy.
@@
Spectral fitting of the combined FXT-A and FXT-B spectra in each region was performed using XSPEC version 12.14.1 over the 0.5–7.0 keV energy range. The diffuse ICM thermal emission was described with the APEC model (version 3.0.9), and the Galactic hydrogen absorption was accounted for using the TBABS model, adopting the Solar abundance table from Anders & Grevesse (1989), with the redshift fixed at 0.0042 and the hydrogen column density fixed at $2.11\times 10^{20}\mathrm{cm}^{-2}$ (Willingale et al. 2013). Other parameters are set to be free. We examined the C-Stat/d.o.f. values for all regions, which were consistently around 1.2. We plotted the spectrum of one region, as shown in Fig. 5. The region belongs to the contour-binning partition and is located in a region approximately $6.6^{\prime}$ from the cluster center. The red and black points represent the spectral data from FXT‐A and FXT‐B in the source region (with the diffuse background subtracted), while the solid line denotes the best‐fit model. The C-Stat/d.o.f. for FXT-A and FXT-B are 246.74/287 and 295.70/299, respectively. The green and blue points display the spectral data from FXT‐A and FXT‐B in the diffuse background region, which is co-located on the CCD with the source region. In addition, the light cyan and orange indicate the particle backgrounds for the source and diffuse background regions, respectively.
@@ 4.1 The map of physical properties
The temperature map (see Fig. 6) reveals a central region in M87 with temperatures below 2.5 keV, along with two distinct low-temperature regions in the northwest and southeast. Sharp temperature increases are seen exterior to these regions, coincident with the locations of two concentric, arc-shaped cold fronts. A low-temperature belt extends outwards from the central region, connecting the low-temperature regions in the southeast and northwest, and spiraling counterclockwise from the northwest, forming a large-scale low-temperature spiral structure. In addition, a region with a lower temperature, compared to the surrounding areas, is located on the outer side of the southwest arm. A detailed analysis is presented in Sect. 4.2.
@@
The metallicity are obtained from the APEC model. The metallicity is lower within the two X-ray arms but higher to the north and south of the center, which is consistent with the findings in Simionescu et al. (2007); Gatuzz et al. (2022). Additionally, beyond $12^{\prime}$, the metallicity in the northwest is significantly higher than that in the southeast, likely due to the accumulation of cold gas in the northwest. In addition, the metallicity of the spiral structure displays higher values, further supporting the existence of the spiral structure.
@@
To investigate its thermodynamic properties, we examine the distributions of pseudo-pressure and pseudo-entropy. The thermal pressure and entropy can be defined (Sasaki et al. 2016) as $P=kTn_{\textnormal{e}}$ and $S=kTn_{\textnormal{e}}^{-2/3}$, respectively, where $n_{e}$ is the electron density and $kT$ is the temperature. Based on the emission measure obtained from spectral fits, we obtained pseudo-entropy and pseudo-pressure. we convert $P$ and $S$ to $P\propto kT{\rm norm}^{0.5}$ and $S\propto kT{\rm norm}^{-1/3}$, respectively (Fabian et al. 2006). Pseudo-pressure and pseudo-entropy maps are also displayed in Fig. 6. In the central region of M87, the pressure map reveals decreasing pressure along the eastern and southwestern X-ray arms, and increasing pressure in the northwest and southeast. The entropy map shows low-entropy features coincident with both X-ray arms, exhibiting an entropy asymmetry towards the northwest and southeast near the core. These results are consistent with those reported in Simionescu et al. (2007). The outer spiral structure also displays lower entropy values, further supporting the existence of the spiral structure.
@@
To sum up, the EP-FXT observations revealed the spiral structure with high completeness and clarity. Moreover, we systematically mapped its thermodynamic properties—including temperature, metal abundance, and entropy—through two-dimensional distributions. These results provide detailed and spatially resolved evidence for the existence and evolution of the spiral structure.
4.2 Sectorial radial profiles
@@
In the residual map (see Fig. 4), we defined four sectorial regions of interest (NW, SE, NE and SW). Using Fig. 1, radial profiles of the surface brightness and temperture are extracted in each region (see Fig. 7). The surface brightness profiles near the cold fronts are fitted with a broken power-law. The best‐fit parameters of the broken power‐law model are listed in Table 3.
@@
In the northeast, two surface brightness drops are detected at $R=7.44\pm 0.04^{\prime}$ and $R=21.40\pm 0.14^{\prime}$, respectively. At $R=21.40\pm 0.14^{\prime}$, a modest increase in temperature is found from $2.16^{+0.11}_{-0.10}\leavevmode\nobreak\ \mathrm{keV}$ to $2.50^{+0.19}_{-0.16}\leavevmode\nobreak\ \mathrm{keV}$ at the $1.7\sigma$ level (The uncertainties are quoted at the $1\sigma$ level). While the temperature change alone is not highly significant, the corresponding surface brightness discontinuity and the location in the outskirts of M87—where the SNR is lower—suggest that this feature likely marks the outer boundary of the northeastern spiral structure.
@@
In the northwest, a drop in surface brightness is found at $R=18.33\pm 0.08^{\prime}$, accompanied by a rise in temperature from $2.39^{+0.06}_{-0.07}$ to $2.70^{+0.12}_{-0.14}\leavevmode\nobreak\ \mathrm{keV}$, corresponding to a significance of approximately $2.2\sigma$. Despite the moderate significance of the temperature jump, the presence of a surface brightness edge, coupled with the low SNR in the outskirts, supports the classification of this feature as a cold front, in agreement with previous findings.
@@
In the southeast, a cold front is identified at $R=6.76\pm 0.03^{\prime}$, indicated by a drop in surface brightness and a temperature increase from $2.43^{+0.03}_{-0.02}$ to $2.58^{+0.04}_{-0.03}\leavevmode\nobreak\ \mathrm{keV}$ at the $3.3\sigma$ level. At $R=17.52\pm 0.22^{\prime}$, the surface brightness shows a mild discontinuity, transitioning from a deficit to an excess in the residual map (see Fig. 4). This feature may correspond to the inner boundary of the southeastern portion of the spiral structure.
@@
A region of excess surface brightness is detected in the southwest at $R=9.75\pm 0.07^{\prime}$, potentially associated with an arc-shaped feature from past AGN activity (Forman et al. 2005). The surface brightness discontinuity is described by a broken power-law model, and a corresponding radial temperature increase further supports the possible presence of a weak cold front.
@@
Table 3: Best-fit broken power-law model parameters.
@@
Figure 7: Radial profiles of surface brightness and temperature in the four sectorial regions marked as NE, NW, SE and SW. The surface brightness were extracted from Fig. 1. Red curves indicates the best-fit broken power-law model. Temperature increases and decreases at density discontinuities are marked by black and green vertical lines, respectively.
@@ 5 Discussion
The imaging and imaging-spectral analyses reveal detailed structures at the center of the Virgo cluster. In addition to the well-known X-ray arms, cavities, and cold fronts, a large-scale sloshing spiral pattern is identified in the GGM and residual images, connecting the two previously recognized cold fronts. Thanks to its low particle background and wide FoV, EP-FXT observations provide a continuous view of the spiral structure in the outskirts of M87, consistent with a similar pattern seen in the mosaic image from XMM-Newton (Simionescu et al. 2010). Building on these observations, we employed EP-FXT data to construct two-dimensional thermodynamic maps of the outskirts of M87, providing further insight into the large-scale physical properties of the spiral structure. In this section, we discuss the properties of the spiral structure and compare the results with those obtained by XMM-Newton.
@@ 5.1 Comparison with XMM-Newton
Figure 8: Comparison of temperature in each pixel between EP-FXT, XMM-Newton. The blue dashed line represents 1:1. The red line represents the best-fit linear regression, shown in the legend.
Using the same contour-binning regions as defined for the EP-FXT data, we performed spectral fitting on the archival XMM-Newton observation (ObsID 0803670601) over the 0.5–7.0 keV energy band, adopting the same model described in Sect. 4. The temperatures for each region are compared with those obtained from EP-FXT, as shown in Fig. 8 and Table LABEL:tab:cmp. In each panel, the blue dashed line denotes the one-to-one correspondence, while the red solid line represents the best-fit linear regression.
As shown in Fig. 8, the temperatures measured by FXT-A and FXT-B are consistent within the uncertainties, although the temperature from FXT-A appears slightly higher than that from FXT-B. This discrepancy may be attributed to the greater effective area of FXT-B at lower energies. Overall, the temperatures obtained from EP-FXT generally agree with those derived from XMM-Newton, despite minor discrepancies. These differences may result from the presence of multi-temperature components within the ICM and calibration differences of the two telescopes, which together can lead to systematic discrepancies in temperature measurements (Migkas et al. 2024; Liu et al. 2023; Schellenberger et al. 2015). Overall, EP-FXT delivers results that are consistent and comparable to those of XMM-Newton.
Table 4: Comparison of temperatures measured by EP-FXT and XMM-Newton.
@@ 5.2 Giant sloshing spiral
Based on the above analysis, we further confirmed the previously identified spiral structure using the GGM and residual images from EP-FXT, which provide a comprehensive view of the feature. Subsequently, we performed imaging-spectral analysis to construct two-dimensional maps of temperature, metallicity, and entropy on the outskirts of M87, providing a spatially resolved characterization of the thermodynamic properties of the spiral structure. The results show that along the extent of the spiral, metallicity is enhanced, while temperature and entropy are reduced. In addition, the spiral structure exhibits significant surface brightness enhancements, with cooler gas preferentially concentrated in the brighter regions, forming an alternating pattern of bright, cold areas and dimmer, hotter regions. Numerical simulations by Roediger et al. (2011) predict that such structures should exhibit low temperatures, high densities, and elevated metallicity. Our observations are consistent with these predictions, supporting the scenario in which the large-scale spiral structure is produced by gas sloshing triggered by a minor merger event (Ascasibar & Markevitch 2006).
Figure 9 shows the large spiral structure in Virgo, detected by EP-FXT, consisting of an inner ring-like structure and outer spiral arms, which are delineated by a white spiral line. In the northeast of the spiral structure (as indicated by the white arrow), the X-ray surface brightness is lower than in the surrounding regions, forming a concave feature. Two additional X-ray concave regions are also present in the west-southwest. This structure is similar to the Kelvin-Helmholtz rolls presented in the simulations and the smeared-out spiral structures observed in A496 (Walker et al. 2017; Ghizzardi et al. 2014). This similarity may be attributed to the influence of the Kelvin-Helmholtz instability (KHI), which generates turbulence and disturbances near the cold front, potentially affecting the formation of the spiral structure.
Overall, EP-FXT provides a clear and comprehensive view of the large-scale spiral pattern in M87. The spiral feature appears to rotate counterclockwise, extending from the cold front in the southeast to the cold front in the northwest, and further spiraling toward the cold gas clump in the southern outskirts. This spiral feature is associated with low temperature, low entropy, and high metallicity, consistent with expectations for gas sloshing phenomena. The prominent and well-defined morphology revealed by EP-FXT offers valuable insights into the formation and evolution of spiral structures in galaxy clusters.
@@ 6 Conclusion
This paper provides a detailed analysis of the images and spectra of the Virgo cluster with EP-FXT. In the images, we present a prominent and complete large-scale spiral structure. Furthermore, we produce two-dimensional maps of temperature, entropy, and metallicity that clearly illustrate the characteristics of the M87 spiral structure. The main results of this work are summarized as follows:
@@
    In the GGM image, multiple structures are identified. Two prominent X-ray arms are clearly visible: the outer side of the southwest arm bends toward the southeast, while the east arm bends toward the northwest. A shock from the northwest appears to connect to the east arm. Furthermore, cold fronts in the southeast and northwest are also readily discernible. These structures are broadly consistent with previous findings.
@@
    Based on the residual map (see Fig. 4), an ring-like structure with enhanced surface brightness is found in the outer X-ray arms. This ring-like structure, connected to the northwestern cold front, rotates counterclockwise, forming a large, outward-extending spiral structure. This structure exhibits low temperature, low entropy, and high density, which are characteristics consistent with gas sloshing induced by the merger of a gasless subcluster.
@@
    Temperatures measured with EP-FXT are consistent with those obtained with XMM-Newton within errors.
    Figure 9: The EP-FXT surface brightness residual image in the 0.3–2.5 keV band shows a spiral structure. White arrowheads indicate locations where this structure is truncated, possibly because of the KHI. The central ring structure of M87 and the outer spiral structure appear to be part of a single, larger structure, as delineated by the white spiral line.

### v3

@@ 4 Imaging spectral analysis
The X-ray image obtained with EP-FXT was divided into 142 regions using the contour-binning software (Sanders 2006) with a S/N of 300. We adopted a flat-weighting assumption, assuming that the X-ray emission within each region is spatially uniform, to calculate the ancillary response files (ARFs). The ARFs were generated using the EP-FXT response tool fxtarfgen. During the calculation, the parameter extend was set to one to account for the extended nature of the source, thus more accurately capturing the spatial distribution of the emission within each extraction region. The spectra were grouped using grppha with a minimum of ten counts per bin to ensure the applicability of the C-statistic. To minimize the impact of background contamination on the source spectra and improve the accuracy of the spectral analysis, we selected observations of the M87 source and its corresponding background, ensuring similar observation times, the same filter settings, and longer exposure durations; specifically, ObsID 13600005124 and ObsID 13600005125. In the spectral analysis, both the particle background and the diffuse background were taken into account. We calculated the particle background for both observations, with a ratio of $1.05\pm 0.03$, indicating that the particle background levels are nearly identical. Therefore, we selected the spectra from the same CCD region in ObsID 13600005125 to serve as the background for ObsID 13600005124. For regions in ObsID 13600005125 that were contaminated by bright point sources or extended sources, background data from nearby uncontaminated areas were used instead. Because our analysis focuses on M87, we intentionally retained the Virgo cluster component in the background data because it reflects the diffuse emission environment surrounding the galaxy better.
@@
We spectrally fit the combined FXT-A and FXT-B spectra in each region using XSPEC version 12.14.1 over the 0.5–7.0 keV energy range. The diffuse ICM thermal emission was described with the APEC model (version 3.0.9), and the Galactic hydrogen absorption was accounted for using the TBABS model, adopting the solar abundance table from Anders & Grevesse (1989), with the redshift fixed at 0.0042 and the hydrogen column density fixed at $2.11\times 10^{20}\mathrm{cm}^{-2}$ (Willingale et al. 2013). Other parameters were set to be free. We examined the C-Stat/d.o.f. values for all regions, which were consistently around 1.2. We plot the spectrum of one region in Fig. 5. The region belongs to the contour-binning partition and is located in a region approximately $6.6^{\prime}$ away from the cluster center. The red and black points represent the spectral data from FXT‐A and FXT‐B in the source region (with the diffuse background subtracted), while the solid line denotes the best‐fit model. The C-Stat/d.o.f. for FXT-A and FXT-B are 246.74/287 and 295.70/299, respectively. The green and blue points display the spectral data from FXT‐A and FXT‐B in the diffuse background region, which is co-located on the CCD with the source region. In addition, the light cyan and orange indicate the particle backgrounds for the source and diffuse background regions, respectively.
@@ 4.1 The map of physical properties
The temperature map (see Fig. 6) reveals a central region in M87 with temperatures below 2.5 keV, along with two distinct low-temperature regions in the northwest and southeast. Sharp temperature increases are seen outside these regions that are coincident with the locations of two concentric arc-shaped cold fronts. A low-temperature belt extends outward from the central region. It connects the low-temperature regions in the southeast and northwest and spirals in counterclockwise from the northwest, forming a large-scale low-temperature spiral structure. In addition, a region with a lower temperature than the surrounding areas is located outside of the southwest arm. A detailed analysis is presented in Sect. 4.2.
@@
The metallicity was obtained from the APEC model. The metallicity is lower within the two X-ray arms, but higher north and south of the center, which is consistent with the findings by Simionescu et al. (2007); Gatuzz et al. (2022). Additionally, beyond $12^{\prime}$, the metallicity in the northwest is significantly higher than that in the southeast, likely due to the accumulation of cold gas in the northwest. In addition, the metallicity of the spiral structure displays higher values, which further supports the existence of the spiral structure.
@@
To investigate its thermodynamic properties, we examined the distributions of pseudo-pressure and pseudo-entropy. The thermal pressure and entropy can be defined (Sasaki et al. 2016) as $P=kTn_{\textnormal{e}}$ and $S=kTn_{\textnormal{e}}^{-2/3}$, respectively, where $n_{e}$ is the electron density, and $kT$ is the temperature. Based on the emission measure obtained from spectral fits, we obtained the pseudo-entropy and pseudo-pressure. We convert $P$ and $S$ into $P\propto kT{\rm norm}^{0.5}$ and $S\propto kT{\rm norm}^{-1/3}$, respectively (Fabian et al. 2006). Pseudo-pressure and pseudo-entropy maps are also displayed in Fig. 6. In the central region of M87, the pressure map reveals decreasing pressure along the eastern and southwestern X-ray arms and increasing pressure in the northwest and southeast. The entropy map shows low-entropy features coincident with both X-ray arms, exhibiting an entropy asymmetry northwest and southeast near the core. These results are consistent with those reported by Simionescu et al. (2007). The outer spiral structure also displays lower entropy values, which further supports the existence of the spiral structure.
@@
To summarize, the EP-FXT observations revealed the spiral structure with high completeness and clarity. Moreover, we systematically mapped its thermodynamic properties, including temperature, metal abundance, and entropy, through two-dimensional distributions. These results provide detailed and spatially resolved evidence for the existence and evolution of the spiral structure.
4.2 Radial sector profiles
@@
In the residual map (see Fig. 4), we defined four sector regions of interest (NW, SE, NE, and SW). Using Fig. 1, we extracted the radial profiles of the surface brightness and temperature in each region (see Fig. 7). The surface brightness profiles near the cold fronts were fit with a broken power law. The best‐fit parameters of the broken power‐law model are listed in Table 3.
@@
In the northeast, two surface brightness drops are detected at $R=7.44\pm 0.04^{\prime}$ and $R=21.40\pm 0.14^{\prime}$. At $R=21.40\pm 0.14^{\prime}$, a modest increase in the temperature is found from $2.16^{+0.11}_{-0.10}\leavevmode\nobreak\ \mathrm{keV}$ to $2.50^{+0.19}_{-0.16}\leavevmode\nobreak\ \mathrm{keV}$ at the $1.7\sigma$ level (the uncertainties are quoted at the $1\sigma$ level). While the temperature change alone is not highly significant, the corresponding surface brightness discontinuity and the location in the outskirts of M87 (where the S/N is lower) suggest that this feature likely marks the outer boundary of the northeastern spiral structure.
@@
In the northwest, a drop in the surface brightness is found at $R=18.33\pm 0.08^{\prime}$, accompanied by a rise in temperature from $2.39^{+0.06}_{-0.07}$ to $2.70^{+0.12}_{-0.14}\leavevmode\nobreak\ \mathrm{keV}$, corresponding to a significance of approximately $2.2\sigma$. Despite the moderate significance of the temperature jump, the surface brightness edge, coupled with the low S/N in the outskirts, supports the classification of this feature as a cold front. This agrees with previous findings.
@@
In the southeast, a cold front is identified at $R=6.76\pm 0.03^{\prime}$, indicated by a drop in the surface brightness and a temperature increase from $2.43^{+0.03}_{-0.02}$ to $2.58^{+0.04}_{-0.03}\leavevmode\nobreak\ \mathrm{keV}$ at the $3.3\sigma$ level. At $R=17.52\pm 0.22^{\prime}$, the surface brightness shows a mild discontinuity, transitioning from a deficit to an excess in the residual map (see Fig. 4). This feature may correspond to the inner boundary of the southeastern portion of the spiral structure.
@@
A region of excess surface brightness is detected in the southwest at $R=9.75\pm 0.07^{\prime}$, which might be associated with an arc-shaped feature from past AGN activity (Forman et al. 2005). The surface brightness discontinuity is described by a broken power-law model, and a corresponding radial temperature increase further supports the possible presence of a weak cold front.
@@
Table 3: Parameters of the best-fit broken power-law model.
@@
Figure 7: Radial profiles of surface brightness and temperature in the four sectorial regions marked NE, NW, SE, and SW. The surface brightness was extracted from Fig. 1. The red curves indicate the best-fit broken power-law model. The temperature increases and decreases at density discontinuities are marked by vertical black and green lines, respectively.
@@ 5 Discussion
The imaging and imaging-spectral analyses revealed detailed structures at the center of the Virgo cluster. In addition to the well-known X-ray arms, cavities, and cold fronts, a large-scale sloshing spiral pattern is identified in the GGM and residual images, where it connects the two previously recognized cold fronts. The low particle background and wide FoV allow the EP-FXT observations to provide a continuous view of the spiral structure in the outskirts of M87. This is consistent with a similar pattern seen in the mosaic image from XMM-Newton (Simionescu et al. 2010). Building on these observations, we employed EP-FXT data to construct two-dimensional thermodynamic maps of the outskirts of M87 to provide further insight into the large-scale physical properties of the spiral structure. In this section, we discuss the properties of the spiral structure and compare the results with those obtained by XMM-Newton.
@@ 5.1 Comparison with XMM-Newton
Figure 8: Comparison of the temperature in each pixel between EP-FXT and XMM-Newton. The dashed blue line represents the 1:1 ratio. The red line represents the best-fit linear regression, shown in the legend.
Using the same contour-binning regions as defined for the EP-FXT data, we performed spectral fitting on the archival XMM-Newton observation (ObsID 0803670601) over the 0.5–7.0 keV energy band, adopting the same model described in Sect. 4. The temperatures for each region were compared with those obtained from EP-FXT, as shown in Fig. 8 and Table LABEL:tab:cmp. In each panel, the blue dashed line denotes the one-to-one correspondence, while the red solid line represents the best-fit linear regression.
As shown in Fig. 8, the temperatures measured by FXT-A and FXT-B are consistent within the uncertainties, although the temperature from FXT-A appears to be slightly higher than that from FXT-B. This discrepancy may be attributed to the greater effective area of FXT-B at lower energies. Overall, the temperatures obtained from EP-FXT generally agree with those derived from XMM-Newton, despite minor discrepancies. These differences may result from the multitemperature components within the ICM and from calibration differences of the two telescopes, which together can lead to systematic discrepancies in the temperature measurements (Migkas et al. 2024; Liu et al. 2023; Schellenberger et al. 2015). Overall, EP-FXT delivers results that are consistent with and comparable to those of XMM-Newton.
Table 4: Comparison of the temperatures measured by EP-FXT and XMM-Newton.
@@ 5.2 Giant sloshing spiral
Based on the above analysis, we further confirmed the previously identified spiral structure using the GGM and residual images from EP-FXT, which provide a comprehensive view of the feature. Subsequently, we performed an imaging-spectrum analysis to construct two-dimensional maps of the temperature, metallicity, and entropy in the outskirts of M87, providing a spatially resolved characterization of the thermodynamic properties of the spiral structure. The results show that throughout the spiral, the metallicity is enhanced, while temperature and entropy are reduced. In addition, the spiral structure exhibits significant surface brightness enhancements. Cooler gas is preferentially concentrated in the brighter regions, where it forms an alternating pattern of bright cold areas and dimmer hotter regions. Numerical simulations by Roediger et al. (2011) predicted that these structures should exhibit low temperatures, high densities, and an elevated metallicity. Our observations are consistent with these predictions and support the scenario in which the large-scale spiral structure is produced by gas sloshing that was triggered by a minor merger event (Ascasibar & Markevitch 2006).
Figure 9 shows the large spiral structure in Virgo as detected by EP-FXT. It consists of an inner ring-like structure and outer spiral arms. Northeast of the spiral structure (as indicated by the white arrow), the X-ray surface brightness is lower than in the surrounding regions and forms a concave feature. Two additional X-ray concave regions are also present in the west-southwest. This structure is similar to the Kelvin-Helmholtz rolls presented in the simulations and the smeared-out spiral structures observed in A496 (Walker et al. 2017; Ghizzardi et al. 2014). This similarity may be attributed to the influence of the Kelvin-Helmholtz instability, which generates turbulence and disturbances near the cold front that might affect the formation of the spiral structure.
Overall, EP-FXT provides a clear and comprehensive view of the large-scale spiral pattern in M87. The spiral feature appears to rotate counterclockwise and extends from the cold front in the southeast to the cold front in the northwest. It further spirals toward the cold gas clump in the southern outskirts. This spiral feature is associated with a low temperature, low entropy, and high metallicity, consistent with expectations for gas-sloshing phenomena. The prominent and well-defined morphology revealed by EP-FXT offers valuable insights into the formation and evolution of spiral structures in galaxy clusters.
@@ 6 Conclusion
We provided a detailed analysis of the images and spectra of the Virgo cluster with EP-FXT. The images show a prominent and complete large-scale spiral structure. Furthermore, we produced two-dimensional maps of the temperature, entropy, and metallicity that clearly illustrated the characteristics of the M87 spiral structure. The main results of this work are summarized below.
@@
    In the GGM image, multiple structures are identified. Two prominent X-ray arms are clearly visible: The outer side of the southwest arm bends toward the southeast, and the east arm bends toward the northwest. A shock from the northwest appears to be connected to the east arm. Furthermore, cold fronts in the southeast and northwest are also readily discernible. These structures are broadly consistent with previous findings.
@@
    Based on the residual map (see Fig. 4), a ring-like structure with an enhanced surface brightness is found in the outer X-ray arms. This ring-like structure is connected to the northwestern cold front and rotates counterclockwise. It forms a large, outward-extending spiral structure. This structure has a low temperature, low entropy, and high density. These characteristics are consistent with gas sloshing induced by the merger of a gasless subcluster.
@@
    The temperatures measured with EP-FXT are consistent with those obtained with XMM-Newton within the errors.
    Figure 9: EP-FXT surface brightness residual image in the 0.3–2.5 keV band showing a spiral structure. The white arrowheads indicate locations where this structure is truncated, possibly because of the Kelvin-Helmholtz instability. The central ring structure of M87 and the outer spiral structure appear to be part of a single larger structure, as delineated by the white spiral line.